OpenStudy (anonymous):

this is a physics question that requires a lot of math. Hoping on of you genius could help me. Q: You're driving down the highway late one night at 16m/s when a deer steps onto the road 44m in front of you. Your reaction time before stepping on the brakes is 0.50s , and the maximum deceleration of your car is 10m/s2 . How much distance is between you and the deer. I solved this =23m What is the maximum speed you could have and still not hit the deer? The answer is 25 m/s. A lot of people had stated to use the quadratic formula. I just don't know where to begin.

OpenStudy (tkhunny):

Have you considered hitting it one piece at a time? Reaction Distance: Speed * 0.50 s Initial Problem: 16 m/s * 0.50 s = 8 m Braking Acceleration: a(t) = -10 m/s^2 -- This is constant. Braking Velocity: s(t) = -10 m/s^2 (t) + "Initial Velocity" Initial Problem: s(t) = -10 m/s^2 (t) + 16 m/s Braking Location: x(t) = -5 m/s^2 (t^2) + 16 m/s (t) - "Initial Location" Initial Problem: x(t) = -5 m/s^2 (t^2) + 16 m/s (t) - (44 m - 8 m) Having said all that, the first question makes no sense. Does it mean the distance when the vehicle stops? s(t) = -10 m/s^2 (t) + 16 m/s = 0 ==> t = 16/10 s = 8/5 s That's how long it takes to stop. Where are we when that happens? x(8/5) = -5 m/s^2 ((8/5 s)^2) + 16 m/s (8/5 s) - (44 m - 8 m) = -23.2 m This seems to be a little different from your response. Now, the challenge is to do this all over again, but missing the initial velocity. Initial Velocity: V -- Let's just call it this so we can talk about it. Initial Distance is the same: 44 m -- No change. Reaction Distance: V/2 Braking Acceleration: a(t) = -10 m/s^2 -- No Change Braking Velocity: s(t) = -10 m/s^2 (t) + V -- Missing V! Braking Location: x(t) = -5 m/s^2 (t^2) + V (t) - (44 m - V/2 m) When do we stop? s(t) = -10 m/s^2 (t) + V = 0 ==> t = V/10 s Where do we stop? x(t) = -5 m/s^2 ((V/10 s)^2) + V (V/10 s) - (44 m - V/2 m) = 0 -- We might just brush the poor, frightened deer. One piece at a time! Slowly. Methodically.

OpenStudy (tkhunny):

BTW: I get 25.083 m/s